OpenStudy (anonymous):

The hypotenuse of a right triangle has one end at the origin and one end on the curve y =4th root of (x)

OpenStudy (anonymous):

You could get lots of answer since the orientation of the legs isn't specified. (You can always draw that segment; and put a 45-45-90 right triangle on it, for example.) If one leg is on the +x axis, though, you can get an answer. The area of a right triangle is half the product of the legs. Pick a point (x,y) on the curve, and the legs are x and y, making A = xy/2. Since y = x^(1/4), you can substitute to get: A = (1/2) x^(5/4)

OpenStudy (anonymous):

Here is a picture of the graph. Sorry!!

OpenStudy (anonymous):

Thanks! That helped a lot! I actually understand this problem now lol!!

OpenStudy (anonymous):

No problem(: